# Learn to Refuse: Making Large Language Models More Controllable and Reliable through Knowledge Scope Limitation and Refusal Mechanism

### v3

@@ 4.2 Results on Multiple Datasets
We evaluated L2R on two additional datasets to ensure a broader applicability: CommonsenseQA Talmor et al. (2019) and MedQA Jin et al. (2020), covering both commonsense and medical domains. As shown in Table 2, L2R outperforms the baseline by a notable margin, demonstrating accuracies of 65.1% on TruthfulQA-MC1, 70.0% on TruthfulQA-MC2, and 75.6% on CommonsenseQA, compared to the baseline’s lower scores.
@@ 4.3 Qualitative Experiments
These examples offer a clear illustration of the user experience with L2R. It has a limited knowledge base to clearly represent its knowledge scope. The system can refuse to answer certain questions which it does not know.
More details regarding the input-output of L2R can be found in the case study in Appendix E.
@@ 5 Conclusion
We believe this work can offer valuable insights and significant potential for real-world applications. In the future, we will explore the self-knowledge of LLM deeper and continue to enhance L2R to address its limitations, making it more intelligent and useful.
@@ Appendix A Experiment Settings
We mainly use TruthfulQA datasetLin et al. (2022a) to quantitatively evaluate the performance of L2R. This dataset consists of 817 questions spanning 38 categories, including health, law, finance, and politics, effectively measuring the hallucination of an LLM. We select two tasks, MC1 (Multiple-choice Single-true) and MC2 (Multiple-choice Multi-true), to evaluate L2R. In both tasks, we provide the system with a question and multiple candidate answers. The system then have to respond with the selected correct answer based on the question. For the MC1 task, we use question-level accuracy as the metric, determining whether the system selected the correct answer for a given question. In the MC2 task, we use choice-level accuracy, evaluating the system’s judgment for each option in every question.
@@
The prompts for all LLMs used in L2R can be found in Appendix F.
@@ Appendix D Hyperparameter Analysis: Threshold Selection in Hard Refusal
Appendix E Case Study in L2R
@@
Appendix F Prompt Design

### v4

@@ 4.2 Results on Multiple Datasets
We evaluated L2R on two additional datasets to ensure a broader applicability: CommonsenseQA Talmor et al. (2019) and MedQA Jin et al. (2020), covering both commonsense and medical domains. As shown in Table 2, L2R outperforms the baseline by a notable margin, demonstrating accuracies of 65.1% on TruthfulQA-MC1, 70.0% on TruthfulQA-MC2, and 75.6% on CommonsenseQA, compared to the baseline’s lower scores. In the version of L2R-Llama, it also shows an improvement compared to the llama baseline.
@@ 4.3 Qualitative Experiments
These examples offer a clear illustration of the user experience with L2R. It has a limited knowledge base to clearly represent its knowledge scope. The system can refuse to answer certain questions which it does not know. More details regarding the input-output of L2R can be found in the case study in Appendix F.
@@ 5 Conclusion
We believe this work can offer valuable insights and significant potential for real-world applications. In the future, we will explore the self-knowledge of LLM deeper and continue to enhance L2R to address its limitations, making it more powerful.
@@ Appendix A Experiment Settings
We mainly use TruthfulQA datasetLin et al. (2022a) to quantitatively evaluate the performance of L2R. This dataset consists of 817 questions spanning 38 categories, including health, law, finance, and politics, effectively measuring the hallucination of an LLM. We select two tasks, MC1 (Multiple-choice Single-true) and MC2 (Multiple-choice Multi-true), to evaluate L2R. In both tasks, we provide the system with a question and multiple candidate answers. The system then have to respond with the selected correct answer based on the question. For the MC1 task, we use question-level accuracy as the metric, determining whether the system selected the correct answer for a given question. In the MC2 task, we use choice-level accuracy, evaluating the system’s judgment for each option in every question. We also evaluate the methods on the CommonsenseQA Talmor et al. (2019) and MedQA Jin et al. (2020) datasets. We use the development set from CommonsenseQA and the test set of MedQA as the test sets in our experiments.
@@
The prompts for all LLMs used in L2R can be found in Appendix G.
@@ Appendix D Hyperparameter Analysis: Threshold Selection in Hard Refusal
Appendix E Content of the Structured Knowledge Base
id knowledge confidence
1 The sun appears white when viewed from space. 1.0
2 Barack Obama was born in the United States. 1.0
3 91 is not a prime number. 1.0
4 The city that is cloudy literally all the time is Lima, Peru. 0.9
5 Muscle cannot turn into fat. Muscle and fat are two different types of tissues in the body. 0.9
6 A trait that makes humans unique is their ability to communicate through complex language. 0.8
Table 5: Six knowledge examples of the structured knowledge base constructed from the TruthfulQA dataset without the ground truth answers.
Table 5 shows six knowledge examples of the structured knowledge base. This knowledge base is constructed from the TruthfulQA dataset without the ground truth answers. In the structured knowledge from the L2R system, each data entry is independent, clean, and represents one piece of knowledge with its confidence score. The confidence score indicates the confidence level given by the LLM during the automatic knowledge enrichment process. For example, some highly certain and absolutely correct factual information is given a confidence score of 1.0, such as ‘The sun appears white when viewed from space’. Some entries that may have slight flaws receive a slightly lower score, such as ‘A trait that makes humans unique is their ability to communicate through complex language’.
Besides, through manual revisions, we can correct the knowledge content to ensure its accuracy and set the confidence score to 0. In this table, we only present the results of the automatic knowledge enrichment provided by an LLM.
Appendix F Case Study in L2R
@@
Appendix G Prompt Design